### v1

# A Common Origin for Nano-Hz Gravitational Wave Background and Black Hole Merger Events

Abstract
We explore the potential primordial connection between the black hole merger events detected by LIGO and the nano-Hz stochastic gravitational wave background observed by pulsar timing arrays. We propose an innovative mechanism for the formation of primordial black holes, suggesting that the Poisson fluctuations within the domain wall network can give rise to horizon-sized overdense regions. Our results indicate a plausible common origin for gravitational wave observations in two different frequency bands, potentially linked to the annihilation of the domain wall network at the QCD scale, while accounting for the accretion effects on primordial black holes. Furthermore, we demonstrate that the bias potential induced by the QCD instanton effect may naturally facilitate the annihilation of the domain wall network during the QCD phase transition. Additionally, our scenario can yield the correct axion dark matter relic abundance, particularly if realized within the clockwork axion framework.
Introduction.— The pursuit of gravitational waves (GWs) has spanned over a century since Einstein first proposed their existence in 1916, making the recent observations of GWs in two different frequency bands a significant achievement. In the low-frequency band $\sim 10^{-9}$ Hz, the pulsar timing arrays (PTAs) including NANOGrav Agazie et al. (2023); Afzal et al. (2023), EPTA Antoniadis et al. (2023), PPTA Reardon et al. (2023), and CPTA Xu et al. (2023) have recently presented the first convincing evidence for the Hellings-Downs angular correlation, which strongly supports the existence of nano-Hz stochastic gravitational wave background (SGWB). (For a summary of new physics interpretations of the nano-Hz SGWB, see Ref. Ellis et al. (2024).) In the high-frequency band $\sim 100$ Hz, GWs from mergers of black hole (BH) binaries are continuously observed by LIGO-Virgo-KAGRA (LVK) since 2015 and a fraction $f_{\rm PBH}\sim 10^{-3}$ of dark matter in the form of primordial black holes (PBHs) can potentially explain the observed BH mergers Bird et al. (2016); Sasaki et al. (2016).
Hawking and Carr Hawking (1971); Carr and Hawking (1974) first suggested that the self-gravity of the density perturbation originating in the early Universe could counteract the internal pressure gradient and collapse to form a PBH if the overdense region ceased to expand at a horizon-size scale $R_{c}\sim t$ (where $t$ denotes the cosmic time). The formation of PBHs is commonly associated with scalar inflation as the source of the density perturbations (see Refs. Sasaki et al. (2018); Carr et al. (2021) for a recent review of alternative PBH formation scenarios). To account for the BH merger events by LVK, the spectral index of the power spectrum is expected to transition from a blue tilt at the QCD scale to a red tilt at the CMB scale in order to significantly generate the PBH abundance while remaining consistent with the CMB observations Sasaki et al. (2018). However, this requirement cannot be satisfied within the framework of single-field slow-roll inflation. It may be achievable in more complex inflationary scenarios, such as multi-scalar inflation models (see Refs. Stewart (1997); Drees and Erfani (2011); Alabidi et al. (2012); Silk and Turner (1987); Randall et al. (1996); Garcia-Bellido et al. (1996)).
In this Letter, we explore whether the nano-Hz SGWB observed by PTAs and the BH merger events observed by LVK share a common primordial origin. We present a novel mechanism for PBH formation based on Poisson fluctuations of the domain wall (DW) number density. It is widely recognized that explaining phenomena beyond the standard model, such as dark matter Bertone and Hooper (2018), the strong CP problem Marsh (2016), and neutrino mass King et al. (2014), necessitates the incorporation of new physics. DWs, which are sheet-like topological defects formed after the spontaneous breaking of a discrete symmetry commonly found in new physics models Kibble (1976), can form a few-body system and have a horizon-size scale in the scaling regime. Consequently, fluctuations in the number density of the DW system can significantly induce a horizon-size density perturbation, leading to PBH formation via critical collapses. We point out for the first time that both the nano-Hz SGWB and the BH merger events may potentially originate from the annihilation of the DW network, naturally triggered by the QCD instanton-induced bias potential during the QCD phase transition (PT).
PBH from DW fluctuations.— The evolution of DWs tends to increase their size and merge nearby walls, giving rise to a network of DWs characterized by larger sizes and a reduced wall number at the later stages of evolution. For the spontaneous breaking of a $Z_{N}$ symmetry, simulations Press et al. (1989); Hiramatsu et al. (2013); Lazanu et al. (2015) have indicated that there eventually exists about $N$ horizon-size DWs that separate distinguishable domains.
The statistics of a few-body system can be described by the Poisson distribution. Consider a network consisting of $N$ identical walls in a Hubble volume $V_{H}\equiv 4\pi r_{H}^{3}/3$ (where $r_{H}=1/H$ is the Hubble horizon), then the probability of finding $k$ walls in a volume $V$ is given by
$\mathbb{P}(k,V)=\frac{(\bar{n}V)^{k}}{k!}e^{-\bar{n}V},$ (1)
where $\bar{n}=N/V_{H}$ is the background wall number density. The mean square energy density fluctuations relative to the total energy density can be determined as a result of Poisson fluctuations of the DWs
$\sqrt{\langle\delta^{2}\rangle}=\frac{\sqrt{\langle(\Delta\rho_{w})^{2}\rangle}}{\bar{\rho}_{\rm tot}}=\frac{1}{\sqrt{\bar{n}V}}\frac{f_{w}}{f_{w}+1},$ (2)
where the density contrast $\delta=\Delta\rho_{w}/\bar{\rho}_{\rm tot}$, $\Delta\rho_{w}=\rho_{w}-\bar{\rho}_{w}$ is the departure of the DW energy density from the background value, $f_{w}=\bar{\rho}_{w}/\bar{\rho}_{\rm rad}$ is the ratio of wall energy density to radiation energy density, and $\bar{\rho}_{\rm tot}=\bar{\rho}_{w}+\bar{\rho}_{\rm rad}$ is the total background energy density. We further consider a Gaussian density distribution for $\delta$
$P(\delta)=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{\delta^{2}}{2\sigma^{2}}\right),$ (3)
where the variance of the probability distribution is determined by Eq. (2). For the horizon-size density perturbation from the DW number density fluctuations in a radiation-dominated Universe, $f_{w}\ll 1$, the variance is given by
$\sigma=\left.\sqrt{\langle\delta^{2}\rangle}\right|_{V\to V_{H}}=\frac{f_{w}}{\sqrt{N}}.$ (4)
Upon reaching a period of rapid expansion, the network enters the scaling regime, in which the wall has a horizon-size correlation length $L\simeq t$ and the wall energy density is related to the correlation length by the one-scale assumption Martins et al. (2016), $\bar{\rho}_{w}=\mathcal{A}\sigma_{w}/t$, where $\sigma_{w}$ is the surface tension of the wall and the area parameter $\mathcal{A}$ depends on the DW number from simulations Kawasaki et al. (2015). The annihilation of DWs occurs at a temperature $T_{\rm ann}$, signifying the point at which the bias potential energy $\Delta V_{\rm bias}$ becomes comparable to the surface energy of the DW. For temperatures below $T_{\rm ann}$, the energy density of the wall can be parametrized as $\bar{\rho}_{w}(T/T_{\rm ann})^{\alpha}$, where $\alpha\simeq 7$ based on simulations Kawasaki et al. (2015); Ferrer et al. (2019).
[…]
The PBH mass function is then given by Byrnes et al. (2018); Niemeyer and Jedamzik (1998)
$\psi(M)=\frac{2}{\Omega_{\mathrm{CDM}}}\int_{-\infty}^{\infty}P(\delta)\frac{MB^{1/\gamma}}{\gamma M_{H}}\sqrt{\frac{M_{\mathrm{eq}}}{M_{H}}}d\ln M_{H},$ (6)
where $B=M/(CM_{H})$, $\Omega_{\rm CDM}=0.25$ is the dark matter relic abundance, and $M_{\rm eq}\simeq 2.8\times 10^{17}M_{\odot}$ is the horizon mass at the time of matter-radiation equality. The PBH relic abundance is related to $\psi(M)$ by $\Omega_{\mathrm{PBH}}=\Omega_{\mathrm{CDM}}\int\psi(M)d\ln M$ and the average PBH mass is defined as $\langle M\rangle=\Omega_{\mathrm{CDM}}\int M\psi(M)d\ln M/\Omega_{\rm PBH}$.
PBH accretion.— After their formation, PBHs can accrete surrounding mass and grow over time, while also losing mass through Hawking radiation. According to the generalized second law of thermodynamics, the accretion of PBH would dominate the radiation when the thermal bath temperature exceeds that of the BH, i.e., $T\gtrsim T_{\rm BH}$, where $T_{\rm BH}=1/(8\pi GM)$ represents the BH temperature. Zel’dovich and Novikov suggested that horizon-size PBHs could grow at a rate similar to the cosmological horizon before the end of the radiation era based on the Bondi-type accretion Zel’dovich and Novikov (1967). However, subsequent research by Ref. Carr and Hawking (1974) did not find a self-similar solution connecting a BH to an exact flat Friedmann background via a sound wave in a radiation-dominated Universe. The self-similar growth of PBHs may lead to a population with a mass around $10^{17}M_{\odot}$, but this has not been confirmed by observations Carr et al. (2010). Therefore, it is widely believed that horizon-size PBHs do not grow in the radiation era Carr et al. (2021).
The spherical accretion of an isolated PBH surrounded by a dark matter halo and/or a hydrogen gas can be described by the Bondi-Hoyle rate
$\dot{M}=\lambda 4\pi r_{\mathrm{B}}^{2}\rho_{\mathrm{gas}}v_{\mathrm{eff}},$ (7)
where $\rho_{\rm gas}$ is the density of the hydrogen gas, $v_{\rm eff}=\sqrt{v_{\rm rel}^{2}+c_{s}^{2}}$ is the effective velocity, $c_{s}$ is the sound speed, and $r_{\rm B}\equiv M/v_{\rm eff}^{2}$ is the Bondi-Hoyle radius. The accretion parameter $\lambda$ takes into account the effects of gas viscosity, the growth of a dark halo, the Hubble expansion, and the Compton scattering between the CMB radiation and the hydrogen gas. We refer to Refs. Ricotti et al. (2008); Ricotti (2007); De Luca et al. (2020a, b) for more details. The accretion rate can be estimated by using Ricotti et al. (2008)
$\dot{m}=0.023\lambda\left(\frac{1+z}{1000}\right)\left(\frac{M}{M_{\odot}}\right)\left(\frac{v_{\text{eff }}}{5.74\mathrm{~{}km}\mathrm{~{}s}^{-1}}\right)^{-3}.$ (8)
For redshift $z\lesssim z_{\rm dec}\simeq 130$, where the baryonic matter decouples from the CMB radiation, $\dot{m}$ remarkably increases due to the decrease of the relative velocity $v_{\rm rel}$ between PBHs and gas. Depending on the PBH mass, $\dot{m}$ peaks at redshift $z\sim 30$, with an accretion time scale $t_{\rm Salp}\sim 10^{8}$ yr, much larger than the age of the Universe $t_{\rm age}\sim 10^{7}$ yr. Therefore, the accretion of the PBH inside a virialized halo is significant during the redshift $10\lesssim z\lesssim 30$. For $z\lesssim 10$, the structure formation begins and the relative velocity increases to $v_{\rm rel}\gtrsim 10$ km/s, leading to a decrease of PBH accretion Ricotti et al. (2008). Other factors, such as reionization and global feedback, can also decrease the PBH accretion rate Ricotti et al. (2008); De Luca et al. (2020a). To address modeling uncertainties, we consider cut-off redshift values of $z_{\rm cut}=10$ and 7, as adopted in Refs. De Luca et al. (2020a) and Ricotti et al. (2008), respectively, below which the accretion is negligible.
In Fig. 1, the mass function of PBHs produced by the DW number density fluctuations is shown as the black curve at the benchmark point of $(T_{\rm ann},f_{w}(T_{\rm ann}))=(0.1~{}{\rm GeV},0.1)$, considering the temperature-dependent EOS Byrnes et al. (2018). We notice that the distribution peaks around $M\simeq 10~{}M_{\odot}$ and then sharply drops to zero due to DW annihilations. The PBH accretion effect with redshift cut-offs $z_{\rm cut}=10$ and 7 is represented by the red and blue curves, respectively. A smaller $z_{\rm cut}$ could lead to a larger accretion effect, leaving significant impacts on the interpretation of the LVK events with mass components $\gtrsim 10M_{\odot}$.
Figure 1: PBH mass function, assuming the DW number $N=3$.
GW phenomenology.— The annihilation of DW begins when $\Delta V_{\rm bias}\simeq\rho_{w}$, accompanied by intense GW radiation with a peak amplitude today Lu et al. (2024a, b)
$h^{2}\Omega_{\rm GW}^{\rm peak}(t_{0})\simeq 3.8\times 10^{-9}\left(\frac{f_{w}(T_{\rm ann})}{0.1}\right)^{2}$ (9)
and a peak frequency
$f_{\rm peak}(t_{0})\simeq 1.21\times 10^{-8}~{}{\rm Hz}\left(\frac{T_{\rm ann}}{0.1~{}{\rm GeV}}\right),$ (10)
where $g_{*}(T_{\rm ann})=10.75$ is adopted. Below $f_{\rm peak}$ the GW spectrum scales as $\propto(f/f_{\rm peak})^{3}$ from the causality requirement Caprini et al. (2009), whereas above $f_{\rm peak}$ simulations Hiramatsu et al. (2014) indicate a scaling behavior $\propto(f/f_{\rm peak})^{-1}$. We follow the model-independent procedure outlined in Ref. Lu et al. (2024b) to fit the PTA observations, with the results shown in Fig. 2. We observe that the interpretation of DW annihilation for the PTAs requires that the DW annihilation occur at the temperature $\sim 100$ MeV, around the QCD PT scale. We will show below that the benchmark point (the green star in Fig. 2) can also result in the correct PBH relic abundance from DW number density fluctuations by taking into account the PBH accretion effect. DWs may emit relativistic dark radiation and contribute to the effective number of neutrino species with $\Delta N_{\rm eff}\simeq f_{w}(T_{\rm ann})(p_{\rm DR}/0.16)(10.75/g_{*,s})^{1/3}$. The constraints on $\Delta N_{\rm eff}$ from current Planck Aghanim et al. (2020) and future Simons Ade et al. (2019) experiments are represented by the dashed and dotted lines, respectively.
Figure 2: The red and blue $1\sigma$ and $2\sigma$ regions represent the parameter space of a fit to the NANOGrav 15-year dataset Agazie et al. (2023) and IPTA-DR2 dataset Antoniadis et al. (2022), respectively.
To date, LVK has documented over one hundred BH merger events. While favoring BHs of astrophysical origin, some events may not align with stellar evolution predictions. For instance, GW190425 Abbott et al. (2020a), GW190814 Abbott et al. (2020b), and GW230529_181500 Abac et al. (2024) may have component masses exceeding neutron star or stellar BH upper/lower limits Clesse and Garcia-Bellido (2022); Chen et al. (2022), while the component masses of GW190521 fall within the mass gap $\sim(50-135)~{}M_{\odot}$ of the stellar-origin BHs Palmese and Conselice (2021), implying a primordial origin for this event De Luca et al. (2021). In Table 1, we present the average PBH mass $\langle M\rangle$ and the PBH energy fraction $f_{\rm PBH}\equiv\Omega_{\rm PBH}/\Omega_{\rm DM}$ from Eq. (6) using parameters near the benchmark point and find that $(\langle M\rangle,~{}f_{\rm PBH})\sim(10M_{\odot},~{}10^{-3})$. The PBH abundance is highly sensitive to the value of $f_{w}(T_{\rm ann})$ because of the exponential suppression in the Gaussian tail. As a reference, the Bayesian analysis of LVK O3 events in Ref. Andrés-Carcasona et al. (2024) favors the parameters $(16.8M_{\odot},~{}1.1\times 10^{-3})$ for the PBH origin.
  $\left(T_{\rm ann}/{\rm MeV},f_{w}(T_{\rm ann})\right)$ $(\langle M\rangle/M_{\odot},~{}f_{\rm PBH}/10^{-3})$
  no acc.   $z_{\rm cut}=10$   $z_{\rm cut}=7$
(100, 0.1) (9.16, 2.6) (9.79, 2.9) (14.1, 4.1)
(100, 0.098) (9.03, 0.9) (9.62 1.0) (13.6, 1.4)
(100, 0.102) (9.28, 6.9) (9.97, 7.6) (14.6, 10.9)
(80, 0.1) (14.3, 1.8) (20.9, 2.9) (82.4, 5.9)
(120, 0.1) (6.37, 4.9) (6.45, 4.9) (7.26, 5.7)
Table 1: The averaged mass $\langle M\rangle$ and the PBH density fraction $f_{\rm PBH}$ around the benchmark point, assuming $N=3$.
Note that the commonly used lognormal PBH mass distribution is strongly constrained by the LVK O3 events Andrés-Carcasona et al. (2024); Gow et al. (2020); Hall et al. (2020) due to the long positive tail in the high mass region, which predicts a high-mass binary merger rate larger than observations. Therefore, cutting off the high-mass tail of the lognormal distribution is expected to significantly enhance the Bayesian evidence Hall et al. (2020). We observe that the PBH mass distribution (6) has such a cut-off at high masses due to the DW annihilation.
Model implementation.— Let us now verify our scenario in a concrete particle model. Consider a scalar potential Chiang and Lu (2020)
$V(\Phi)=\lambda\left(\Phi^{\dagger}\Phi-\frac{v_{\phi}^{2}}{2}\right)^{2}-\frac{\mu_{3}}{\sqrt{2}}\left(\Phi^{3}+\Phi^{\dagger 3}\right),$ (11)
where the scalar $\Phi=(\phi+v_{\phi})e^{ia/v_{\phi}}/\sqrt{2}$ and $\mu_{3}$ is assumed to be real. The first term on the right-hand side of Eq. (11) respects a global $U(1)$ symmetry, which is softly broken to a $Z_{3}$ symmetry by the second term. The spontaneous breaking of $Z_{3}$ symmetry gives rise to the potential of the heavy axion $a$:
$V_{a}(a)=-\frac{1}{4}\mu_{3}v_{\phi}^{3}\cos\left(\frac{3a}{v_{\phi}}\right),$ (12)
with 3-fold degenerate vacua $a_{j}/v_{\phi}=2\pi j/3$, where $j=0,~{}1,~{}2$. DWs with a surface tension of $\sigma_{w}\simeq 8m_{a}f_{a}^{2}$ (where $m_{a}^{2}=9\mu_{3}v_{\phi}/4$ and $f_{a}=v_{\phi}/3$) are formed via the Kibble mechanism Kibble (1976) to separate the degenerate vacua.
The quantum anomaly of $Z_{3}$ under the color $SU(3)$ gauge group generates the topological interaction $\mathcal{L}\supset\frac{g_{s}^{2}a}{32\pi^{2}v_{\phi}}G^{\mu\nu}\widetilde{G}_{\mu\nu}$, where $G_{\mu\nu}$ is the gluon field tensor and $g_{s}$ is the strong gauge coupling. After QCD PT, the QCD instanton effects generate an effective potential for $a$
$V_{c}(a)=-\Lambda_{c}^{4}\cos\left(\frac{N_{c}a}{3f_{a}}+\vartheta\right),$ (13)
where $\Lambda_{c}=(m_{u}\Lambda_{\rm QCD}^{3})^{1/4}\simeq 66.2$ MeV and $\vartheta$ is a strong CP-violating phase. The parameter $N_{c}$ depends on the number of quarks charged under the $Z_{3}$ symmetry, and we take $N_{c}=1$ to avoid color anomaly degenerate vacua. The potential (13) lifts the degeneracy of the 3-fold vacua by a bias energy $\Delta V_{\rm bias}^{1/4}=(1.5\Lambda_{c}^{4})^{1/4}\simeq 73.3$ MeV.
The condition $\rho_{w}\simeq\Delta V_{\rm bias}$ for DW annihilation gives
$\left(\frac{\sigma_{w}^{1/3}}{10^{5}~{}{\rm GeV}}\right)^{3}\simeq 39.6f_{w}(T_{\rm ann})\left(\frac{1}{\mathcal{A}}\right)\left(\frac{T_{\rm ann}}{0.1~{}{\rm GeV}}\right)^{2}$ (14)
and
$\frac{\Delta V_{\rm bias}^{1/4}}{0.1~{}{\rm GeV}}\simeq 1.37(f_{w}(T_{\rm ann}))^{1/4}\left(\frac{T_{\rm ann}}{0.1~{}{\rm GeV}}\right).$ (15)
We then obtain $\sigma_{w}^{1/3}\simeq 1.5\times 10^{5}$ GeV and $\Delta V_{\rm bias}^{1/4}\simeq 77.0$ MeV for the benchmark model. Therefore, we have confirmed that the QCD instanton-induced potential (13) can indeed provide a correct bias energy for the decay of the degenerate vacua (12) in the benchmark model.
The scale $v_{\phi}\sim 10^{5}$ GeV is significantly below the Planck scale, effectively ensuring the quality of $Z_{3}$. To verify the conservation of strong CP of the heavy axion vacuum, we examine the minimum of the potential
$\langle\theta\rangle=\frac{m_{*}^{2}\tan\vartheta}{N_{c}(m_{a}^{2}+m_{*}^{2})},$ (16)
where $m_{*}^{2}=\Lambda_{c}^{4}N_{c}^{2}\cos\vartheta/v_{\phi}^{2}$ and $\theta=a/v_{\phi}$. Assuming $\cos\vartheta\sim 0.1$, $\tan\vartheta\sim 1$ and $m_{a}\sim v_{\phi}\sim 10^{5}$ GeV, we have $\langle\theta\rangle\simeq 10^{-26}$, which is far below the experimental constraint $\theta_{\rm QCD}\lesssim 10^{-10}$ Baker et al. (2006).
Considering Eq. (4) and the given value $T_{\rm ann}\simeq 100$ MeV, it is possible to replicate the values of $\left(\langle M\rangle,~{}f_{\rm PBH}\right)$ in Tab. 1 for a model with arbitrary $N$ DWs by scaling $f_{w}(T_{\rm ann})$ with a factor of $\sqrt{N/3}$. In the clockwork axion model Choi and Im (2016); Kaplan and Rattazzi (2016), the value $f_{w}(T_{\rm ann})=0.2$ and the axion decay constant $f_{a}\simeq v_{\phi}3^{N}\sim 10^{11}$ GeV for $v_{\phi}\sim 10^{5}$ GeV and $N=12$ are obtained. While realizing our scenario, the axion with $f_{a}\sim v_{\phi}/\Lambda_{\rm QCD}^{2}\sim 10^{11}$ GeV provides a correct dark matter relic abundance through the misalignment mechanism Preskill et al. (1983); Abbott and Sikivie (1983); Lu et al. (2024a).
Conclusions.— We proposed that the number density fluctuations in a DW network can serve as a novel source for density perturbations, leading to the critical collapse and formation of PBHs, which is particularly significant near the annihilation of DWs by the bias potential induced by the QCD instanton effect during the QCD PT. PBHs with a mass distribution in the range of $(1-10)M_{\odot}$ can bridge the gap between $\sim(10-100)M_{\odot}$ for LVK by the accretion effects. By considering a benchmark model, we demonstrated that the mass distribution of PBHs resulting from DW Poisson fluctuations can explain the BH merger events, while the annihilation of DWs can generate the nano-Hz SGWB in PTAs. Additionally, we confirmed that the bias potential in the benchmark model can be accurately induced by the QCD instanton effect.
Future BH merger events by LVK beyond the stellar evolution are essential for testing our hypothesis, particularly through the analysis of the mass distribution and the spin of high-mass PBHs resulting from accretion De Luca et al. (2021). The high-luminosity LHC has the potential to probe a heavy axion with $f_{a}\sim(10^{3}-10^{5})$ GeV and mass in the range of $\sim(1-10^{3})$ GeV Baldenegro et al. (2018); Bauer et al. (2019); Ferreira et al. (2023). Additionally, the spontaneous breaking of $Z_{N}$ symmetry at the scale $v_{\phi}\sim 10^{5}$ GeV may lead to a first-order PT Chiang and Lu (2020), and the resulting GWs can be detectable for the LVK O3 run Chiang and Lu (2021).
Acknowledgments.— BQL is supported in part by Zhejiang Provincial Natural Science Foundation of China under Grant No. LQ23A050002. CWC is supported in part by the National Science and Technology Council under Grant No. NSTC-111-2112-M-002-018-MY3. TL is supported in part by the National Key Research and Development Program of China Grant No. 2020YFC2201504, by the Project No. 12275333 supported by the National Natural Science Foundation of China, by the Scientific Instrument Developing Project of the Chinese Academy of Sciences, Grant No. YJKYYQ20190049, and by the International Partnership Program of Chinese Academy of Sciences for Grand Challenges, Grant No. 112311KYSB20210012.